OpenStudy (anonymous):

the dallas mavericks finished with the best record for the 06-07 NBA regular season. they won 7 more than four times as many games as they lost. they played 82 games. how many games did they win and lose?

OpenStudy (anonymous):

x = games won y = games lost Two equations can be derived form this, x = 7 + 4*y x + y = 82 Plug in and solve.

OpenStudy (anonymous):

Sooo, if i do.. x + y = 82 x + 4(4) = 82 -1 + 16 = 82 15 = 82 -15 -15 x=67 wins but how many losses? kinda confused sorry..it seemed simple

OpenStudy (anonymous):

Plug in x = 7 + 4*y into the other equation--> (7 + 4*y) + y = 82 5y = 75 --> y = 15 losses and x = 7 + 4*(15) = 67 wins (you were correct)

OpenStudy (anonymous):

okay awesome